# Finding Phones Fast: Low-Latency and Scalable Monitoring of Cellular Communications in Sensitive Areas

### v2

@@ A.0.2. Comparison of Uplink Message Types
The measurement error of the features depends on the message type and its size. As shown in 9(b), the estimated signal power value based on correlation peaks is more accurate for PRACH and PUSCH with more PRBs than for PUCCH and PUSCH with only 1 PRB. This is because the PRACH includes the longest reference signal, allowing for more accurate correlation computations. In contrast, a PUSCH transmission with 1 PRB has only 2 reference signal sequences, each just 12 samples long, leading to poorer performance. A PUCCH transmission, despite spanning only 1 PRB, contains reference signals in 6 OFDM symbols, resulting in better performance than PUSCH with 1 PRB.
@@ A.0.3. Signal Quality
To estimate signal quality, we use a more robust measurement technique compared to previous uplink sniffers (kotuliak_ltrack_2022, ; hoang_ltesniffer_2023, ). In their work, the circular cross-correlation is smoothed with a filter in the frequency domain to eliminate the impact of noise. The difference between the smoothed and non-smoothed version of the correlation is their resulting noise estimate. We call this technique Smoothed Correlation SNR Estimation. For our estimation technique, called Correlation Peak-to-Average measurement, we divide the peak value of the cross-correlation with the average value of the cross-correlation. The intuition is that the signal gets concentrated in and around the peak, while the noise is spread over the whole correlation. The comparison of the two features can be found in 9(c).

### v3

@@ 8. Conclusion
Acknowledgements.
The authors would like to thank Patrick Leu for valuable discussions regarding the system design and for insightful suggestions, including the concept of using physical walls to increase the effect of opposing-facing antennas, which helped shape the final implementation.
@@ A.0.2. Comparison of Uplink Message Types
The measurement error of the features depends on the message type and its size. As shown in 8(b), the estimated signal power value based on correlation peaks is more accurate for PRACH and PUSCH with more PRBs than for PUCCH and PUSCH with only 1 PRB. This is because the PRACH includes the longest reference signal, allowing for more accurate correlation computations. In contrast, a PUSCH transmission with 1 PRB has only 2 reference signal sequences, each just 12 samples long, leading to poorer performance. A PUCCH transmission, despite spanning only 1 PRB, contains reference signals in 6 OFDM symbols, resulting in better performance than PUSCH with 1 PRB.
@@ A.0.3. Signal Quality
To estimate signal quality, we use a more robust measurement technique compared to previous uplink sniffers (kotuliak_ltrack_2022, ; hoang_ltesniffer_2023, ). In their work, the circular cross-correlation is smoothed with a filter in the frequency domain to eliminate the impact of noise. The difference between the smoothed and non-smoothed version of the correlation is their resulting noise estimate. We call this technique Smoothed Correlation SNR Estimation. For our estimation technique, called Correlation Peak-to-Average measurement, we divide the peak value of the cross-correlation with the average value of the cross-correlation. The intuition is that the signal gets concentrated in and around the peak, while the noise is spread over the whole correlation. The comparison of the two features can be found in 8(c).
@@
Appendix B Robustness Test
Figure 9. Effect of uplink receivers being switched off.
We performed robustness tests to determine how well the model behaves if fewer uplink receivers are available or if fewer messages are available for classification. For the first test, we plot these results in Figure 9. In the second test, with just one message (PRACH), the model’s accuracy is 99.17%, and with the first two messages (PRACH and PUSCH (RRC Connection Request)), the model’s accuracy improves to 99.6%.